# Advances in Nonmonotone Proximal Gradient Methods merely with Local Lipschitz Assumptions in the Presense of Kurdyka–Łojasiewicz Property: A Study of Average and Max Line Search

### v1

@@ 3 NPG with Average Line Search and Convergence Results
In order to guarantee the convergence of Algorithm 3.1, we need the following requirement.
@@
Note that Assumption 3.2 (a) and (b) are the same as [18, Assumption 3.1]. Assumption 3.2 (a) guarantees that the optimization problems (Q) are solvable. Assumption 3.2 (b) implies that $g$ is coercive, which consequently implies that subproblems (3.1) always have solutions, and hence the algorithm is well-defined.
Combination of [18, Lemma 3.1] and [12, Lemma 4.1] implies that Algorithm 3.1 is well-defined. The following results are fundamental for the convergence analysis.
@@
    Every iterate $x^{k}$ remains in the sublevel set $\mathcal{L}_{q}(x^{0})\subset\operatorname{dom}g$.
@@
Note that (a)$\sim$(d) have been observed in [12, Lemma 4.2] and [12, Lemma 4.3], so we omit the proof. It remains to consider (e).
@@
holds for all $k\in\mathcal{K}$. We immediately obtain that $\|\hat{x}^{k}-x^{k}\|\to_{\mathcal{K}}0$ (otherwise, the left-hand side in (3.5) goes to infinity) and consequently $\hat{x}^{k}\to_{\mathcal{K}}\bar{x}$ from Assumption 3.2 (b). By Assumption 3.2 (a), using the mean-value theorem yields the existence of a point $\xi^{k}$ on the line segment connecting $x^{k}$ with $\hat{x}^{k}$ such that
@@
Substituting the resulting expression for $g(\hat{x}^{k})-g(x^{k})$ into (3.5) yields
@@
By means of the proof of Lemma 3.4 and the technique of [15, Lemma 4.1], we immediately have the following desired result, i.e., around the accumulation point $\bar{x}$ of $\{x^{k}\}_{k\in\mathbb{N}}$, the associated stepsize sequence remains (uniformly) bounded. This plays a central role in the convergence of entire sequence. We omit the proof because it is highly similar with [17, Lemma 4.1].
@@
Let Assumption 3.2 hold, $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1, and $\bar{x}$ is an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$. Then, for any $\rho>0$, there is a constant $\bar{\gamma}_{\rho}>0$ (usually depending on $\rho$) such that $\gamma_{k}\leq\bar{\gamma}_{\rho}$ for all $k\in\mathbb{N}$ such that $x^{k}\in B_{\rho}(\bar{x})$.
The following result provides the main global (subsequential) convergence of Algorithm 3.1.
@@
From the fact that $g$ is lower semicontinuous, one then has
@@
Theorem 3.6 actually obtained the subsequential convergence, we next aim to give the convergence of the entire sequence. Let sufficiently small $\eta>0$ be the corresponding constant from the definition of the associated desingularization function $\chi$. In view of Proposition 3.3 (d), one can find a sufficiently large index $\hat{k}\in\mathbb{N}$ such that
@@
Finally, throughout the section, let $L_{\rho}>0$ be a (global) Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11). In view of Lemma 3.5, one has
@@
satisfies $\alpha<\frac{1}{2}$, where $\rho>0$, $\bar{\gamma}_{\rho}>0$, and $l>0$ are the constants defined in (3.10), (3.12), and (3.13), respectively, while $L_{\rho}>0$ is a Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11) and $\delta>0$ as well as $\gamma_{\min}>0$ are the parameters from Algorithm 3.1.
@@
From $x^{k}\to_{\mathcal{K}}\bar{x}$, then there exists a sufficiently large $k_{0}\in\mathcal{K}$ satisfying $\Phi_{k_{0}}-q(\bar{x})\geq 0$ sufficiently small deduced by Proposition 3.3 and Theorem 3.6. Recall that that $q$ satisfies KL property at $\bar{x}$, the corresponding disingularizatioon function is denoted as $\chi$. For such $\chi$, one has $\chi(0)=0$ and $\chi$ is continuous and increasing on its domain, therefore one knows $\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}\geq 0$ sufficiently small. Hence, each summand on the right-hand side (3.15) by taking an index $k_{0}\in\mathcal{K}$ sufficiently large yields $\alpha<\frac{1}{2}$. ∎
Using these notations, we get the following result. We omit the proof, since it is highly similar with [17, Lemma 4.3].
@@
By employing the above results, the following theorem demonstrates that the whole sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 3.1 is convergent to its accumulation point $\bar{x}$, provided that $q$ satisfies the KL property at $\bar{x}$.
@@
By Proposition 3.3 (a) and (c), one knows that the whole sequence $\{\Phi_{k}\}_{k\in\mathbb{N}}$ is monotonically decreasing and converges to $q(\bar{x})$ from Theorem 3.6. This implies that $\Phi_{k}\geq q(\bar{x})$ holds for all $k\in\mathbb{N}$. If $\Phi_{k}=q(\bar{x})$ holds for some index $k\in\mathbb{N}$, which, by monotonicity, implies that $\Phi_{k+1}=q(\bar{x})$. Let us now recall Step 6 in Algorithm 3.1, one has
@@
which says that $x^{k+1}=x^{k}$ holds. Since, by assumption, the subsequence $\{x^{k}\}_{\mathcal{K}}$ converges to $\bar{x}$, this implies that $x^{k}=\bar{x}$ for all $k\in\mathbb{N}$ sufficiently large. In particular, one has convergence of the entire (eventually constant) sequence $\{x^{k}\}_{k\in\mathbb{N}}$ to $\bar{x}$ in this situation.
It remains to consider the case where $\Phi_{k}>q(\bar{x})$ for all $k\in\mathbb{N}$. Let $\alpha$ be the constant from (3.15), and $k_{0}\in\mathcal{K}$ be the corresponding iteration index which is used in the definition of $\alpha$, see Lemma 3.7. One then has $0<\Phi_{k}-q(\bar{x})\leq\Phi_{k_{0}}-q(\bar{x})$ for all $k\geq k_{0}$. Without loss of generality, we may also assume that $k_{0}\geq\hat{k}$ defined by (3.9) and that $k_{0}$ is sufficiently large to satisfy
@@
Let $\chi\colon[0,\eta]\to[0,\infty)$ be the desingularization function which comes along with the validity of the KL property of $q$ at $\bar{x}$. Due to $\chi(0)=0$ and $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$, one obtains
@@
hence (3.19) holds for $k=k_{0}$ due to $11/l>$ 1. Suppose that both statements hold for some $k\geq k_{0}$. Using the triangle inequality, the induction hypothesis, and the definition of $\alpha$, one obtains
@@
In particular, it follows from (a) that $x^{k}\in B_{\alpha}(\bar{x})$ for all $k\geq k_{0}$. And taking $k\to\infty$ in (3.19) therefore shows that $\{x^{k}\}_{k\in\mathbb{N}}$ is a Cauchy sequence and, thus, convergent. Since one already knows that $\bar{x}$ is an accumulation point, it follows that the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$. ∎
Note that if all index $k\geq k_{0}$ belong to $S$ or if $p_{k}:=1$, then Algorithm 3.1 degenerates into monotone proximal gradient, see [18, Algorithm 3.1] or [15, Algorithm 3.1], and the corresponding rate-of-convergence result was obtained in [14]. So, we in the following state the rate-of-convergence result of the nonmonotone proximal gradient methods (in general case), readers may find more details for the proof in [17, Theorem 4.6].
@@
4 NPG with Max Line Search and Convergence Analysis
This section focuses on another type of nonmonotone proximal gradient method, where the line search is chosen as max-type rule. Let us now introduce the algorithm which is exactly from [18, Algorithm 4.1].
Algorithm 4.1 Nonmonotone Proximal Gradient Method with Max Line Search
@@ 4 NPG with Max Line Search and Convergence Analysis
Throughout the section, let us denote $l(k)\in\{k-m_{k},\ldots,k\}$ an index satisfying
@@
Note that Assumption 3.2 (b) and Assumption 4.2 imply that $q$ is uniformly continuous on the corresponding sublevel $\mathcal{L}_{q}(x^{0})$. The requirement of uniformly continuity about objective function plays an essential role in the context of nonmonotone line search rules with maximum taste [13]. We here recall the results again for the convenience of readers.
@@
Let Assumption 3.2 and Assumption 4.2 hold and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1, then
@@
    the sequence $\{q(x^{l(k)})\}$ is monotonically decreasing,
@@
    the sequences $\{x^{k}\}$, $\{x^{l(k)}\}$ $\subset\mathcal{L}_{q}(x^{0})\subset\operatorname{dom}g$,
@@
    suppose $\bar{x}$ is an accumulation point of $\{x^{k}\}$ such that $x^{k}\to_{\mathcal{K}}\bar{x}$ holds. Then $\bar{x}$ is an M-stationary point of (Q), and $\gamma_{k}\|x^{k+1}-x^{k}\|\to_{\mathcal{K}}0$ is valid.
With the aid of Proposition 4.3 (d) as well as the techniques in Lemma 3.5, we can obtain the following results.
@@
Let Assumption 3.2 and Assumption 4.2 hold, let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1, $\bar{x}$ is an accumulation point of $\{x^{k}\}$ and consequently let $\{x^{k}\}_{\mathcal{K}}$ be a subsequence converging to some point $\bar{x}$. Then, for any $\rho>0$, there is a constant $\bar{\gamma}_{\rho}>0$ (usually depending on $\rho$) such that $\gamma_{k}\leq\bar{\gamma}_{\rho}$ for all $k\in\mathbb{N}$ such that $x^{k}\in B_{\rho}(\bar{x})$.
We now discuss the convergence properties of the sequence generated by Algorithm 4.1 in the presence of KL property of $q$ at the mentioned accumulation point $\bar{x}$. Let sufficiently small $\eta>0$ be the corresponding constant from the definition of the associated desingularization function $\chi$. In view of Lemma 4.4, we can find a sufficiently large index $\hat{k}\in\mathbb{N}$ such that
@@
Consequently, let $L_{\rho}>0$ be a (global) Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (4.6). In view of Lemma 4.4, let $\bar{\gamma}_{\rho}>0$ be satisfying
@@
Thanks to the above notation, we can find a new neighborhood centered at the accumulation point, whose radium is very small which is mainly used to to ensure the convergence of the whole sequence under the assumption of KL property of $q$, we now introduce such radium.
@@
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$ with desingularization function $\chi$. Then there is a sufficiently large constant $k_{0}\in\mathcal{K}\cap\overline{K}$ such that $k_{0}-m-1$ is also sufficiently large, where $K$ is defined by (4.4), then the corresponding constant is defined as
@@
$\leq\left(\frac{8m(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4m}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
@@
From the fact that $\{x^{k}\}_{\mathcal{K}}\to\bar{x}$, the decreasement of $\{q(x^{l(k)})\}$, we have $\|x^{k_{0}}-\bar{x}\|$ is sufficiently small by the sufficiently large $k_{0}\in\mathcal{K}$, and $q(x^{l(k)})-q(\bar{x})\to 0$ from Assumption 4.2. Then we have $\chi(q(x^{l(k)})-q(\bar{x}))\to 0$ from the continuity of $\chi$ and $\chi(0)=0$. Meanwhile, because $k_{0}-m-1$ is sufficiently large, we have $q(x^{l(k_{0}-m-1)})-q(\bar{x})$ is sufficiently small. Totally, we have $\alpha<\frac{1}{2}$. ∎
Taking the similar analysis with Lemma 3.8, we obtain the following result.
@@
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, then
@@
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$. If $q(x^{l(k)})\neq q(\bar{x})$ for all $k\in\mathbb{N}$ and for any $v\geq k_{0}$ where $k_{0}$ is used to define $\alpha$ in Lemma 4.5, if $x^{k}\in B_{\alpha}(\bar{x})$ holds for all $k_{0}\leq k\leq v$, then we have
@@
Let $\chi:[0,\eta]\to[0,\infty)$ be the desingularization function which comes along with the validity of the KL property of $q$. Due to $\chi(0)=0$ and $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$, one obtains

### v3

@@ 3 The NPG Method with Average Line Search and Convergence Results
In order to guarantee the convergence of Algorithm 3.1, we introduce the following requirements.
@@
Note that Assumption 3.2 (a) and (b) are the same as those presented in [19, Assumption 3.1]. Assumption 3.2 (a) guarantees that the optimization problem (Q) is solvable. Assumption 3.2 (b) implies that $g$ is coercive, which consequently indicates that subproblems (3.1) always have solutions, and hence the algorithm is well-defined.
The combination of [19, Lemma 3.1] and [13, Lemma 4.1] implies that Algorithm 3.1 is well-defined. The following results are fundamental to the convergence analysis.
@@
    Every iterate $x^{k}$ remains in the sublevel set $\mathcal{L}_{q}(x^{0}):=\{x\in\operatorname{dom}q\,|\,q(x)\leq q(x^{0})\}\subset\operatorname{dom}g$.
@@
Note that (a)--(d) have been observed in [13, Lemma 4.2] and [13, Lemma 4.3], so we omit the proof. It remains to consider (e).
@@
holds for all $k\in\mathcal{K}$. We immediately obtain that $\|\hat{x}^{k}-x^{k}\|\to_{\mathcal{K}}0$ (otherwise, the left-hand side in (3.5) goes to infinity) and consequently $\hat{x}^{k}\to_{\mathcal{K}}\bar{x}$ from Assumption 3.2 (b). By Assumption 3.2 (a), the mean-value theorem yields that there exists $\xi^{k}$ on the line segment connecting $x^{k}$ with $\hat{x}^{k}$ such that
@@
Recall (3.5), we have
@@
By means of the proof of Lemma 3.4 and the technique outlined in [16, Lemma 4.1], we immediately obtain the following desired result, i.e., around the accumulation point $\bar{x}$ of $\{x^{k}\}_{k\in\mathbb{N}}$, the associated stepsize sequence remains (uniformly) bounded. We omit the proof because it closely resembles that of [18, Lemma 4.1].
@@
Let Assumption 3.2 hold, $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 3.1, and $\bar{x}$ be its accumulation point. Then, for any $\rho>0$, there is a constant $\bar{\gamma}_{\rho}>0$ (depending on $\rho$) such that $\gamma_{k}\leq\bar{\gamma}_{\rho}$ for all $k\in\mathbb{N}$ such that $x^{k}\in B_{\rho}(\bar{x})$.
The following result refers to the subsequential convergence of Algorithm 3.1.
@@
From the fact that $g$ is lower semicontinuous, then one has
@@
Theorem 3.6 actually obtained the subsequential convergence, we next aim to give the convergence of the entire sequence in the presence of the KL property. Let sufficiently small $\eta>0$ be the corresponding constant from the definition of the associated desingularization function $\chi$. In view of Proposition 3.3 (d), one can find a sufficiently large index $\hat{k}\in\mathbb{N}$ such that
@@
Finally, let $L_{\rho}>0$ be the Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11). In view of Lemma 3.5, one has
@@
satisfies $\alpha<\frac{1}{2}$, where $\rho>0$, $\bar{\gamma}_{\rho}>0$, and $l>0$ are the constants defined in (3.10), (3.12), and (3.13), respectively, while $L_{\rho}>0$ is the Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (3.11) and $\delta>0$ as well as $\gamma_{\min}>0$ are the parameters from Algorithm 3.1.
@@
From $x^{k}\to_{\mathcal{K}}\bar{x}$, then there exists a sufficiently large $k_{0}\in\mathcal{K}$ satisfying $\Phi_{k_{0}}-q(\bar{x})\geq 0$ sufficiently small deduced by Proposition 3.3 and Theorem 3.6. Recall that that $q$ satisfies the KL property at $\bar{x}$, the corresponding disingularizatioon function is denoted as $\chi$. For such $\chi$, one has $\chi(0)=0$ and $\chi$ is continuous and increasing on its domain, therefore $\chi\big{(}\Phi_{k_{0}}-q(\bar{x})\big{)}\geq 0$ is sufficiently small. Hence, each summand on the right-hand side (3.15) by taking an index $k_{0}\in\mathcal{K}$ sufficiently large yields $\alpha<\frac{1}{2}$. ∎
Using these notations, we obtain the following result. The proof is omitted because it is similar to [18, Lemma 4.3].
@@
By employing the above results, the following theorem demonstrates that the whole sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 3.1 convergences to its accumulation point $\bar{x}$, provided that $q$ satisfies the KL property at $\bar{x}$.
@@
By Proposition 3.3 (a) and (c), one knows that the whole sequence $\{\Phi_{k}\}_{k\in\mathbb{N}}$ is monotonically decreasing and converges to $q(\bar{x})$ from Theorem 3.6. It implies that $\Phi_{k}\geq q(\bar{x})$ holds for all $k\in\mathbb{N}$. If $\Phi_{k}=q(\bar{x})$ holds for some index $k\in\mathbb{N}$, which, by monotonicity, implies that $\Phi_{k+1}=q(\bar{x})$. Let us now recall Step 6 in Algorithm 3.1, one has
@@
which says that $x^{k+1}=x^{k}$ holds. Since the subsequence $\{x^{k}\}_{\mathcal{K}}$ are assumed to converge to $\bar{x}$, this implies that $x^{k}=\bar{x}$ for all $k\in\mathbb{N}$ sufficiently large. Consequently, the entire (eventually constant) sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$ in this situation.
It remains to consider the case where $\Phi_{k}>q(\bar{x})$ for all $k\in\mathbb{N}$. Let $\alpha$ be the constant from (3.15), and $k_{0}\in\mathcal{K}$ be the corresponding iteration index used in the definition of $\alpha$, see Lemma 3.7. One then has $0<\Phi_{k}-q(\bar{x})\leq\Phi_{k_{0}}-q(\bar{x})$ for all $k\geq k_{0}$. Without loss of generality, we may also assume that $k_{0}\geq\hat{k}$ defined by (3.9) and that $k_{0}$ is sufficiently large to satisfy
@@
Let $\chi\colon[0,\eta]\to[0,\infty)$ be the desingularization function which validates the KL property of $q$ at $\bar{x}$. Due to $\chi(0)=0$ and $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$, one obtains
@@
hence (3.19) holds for $k=k_{0}$ due to $11/l>$ 1. Suppose that both statements hold for some $k\geq k_{0}$. By the triangle inequality, the induction hypothesis, and the definition of $\alpha$, one obtains
@@
It follows from (a) that $x^{k}\in B_{\alpha}(\bar{x})$ for all $k\geq k_{0}$. Taking $k\to\infty$ in (3.19) shows that $\{x^{k}\}_{k\in\mathbb{N}}$ is a Cauchy sequence and hence convergent. Since $\bar{x}$ is an accumulation point, then the entire sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$. ∎
Note that if all indices $k\geq k_{0}$ belong to $S$ or if $p_{k}:=1$, then Algorithm 3.1 degenerates into a monotone proximal gradient method, see [19, Algorithm 3.1] or [16, Algorithm 3.1], and the corresponding rate-of-convergence result was obtained in [15]. So, we in the following state the rate-of-convergence result of the NPG method (in the general case). Readers may find more details for the proof in [18, Theorem 4.6].
@@
4 The NPG method with Max Line Search and Convergence Analysis
This section focuses on another type of the nonmonotone proximal gradient method, where the line search is chosen as the max-type rule. Let us now introduce the algorithm which is derived from [19, Algorithm 4.1].
Algorithm 4.1 A Nonmonotone Proximal Gradient Method with Max Line Search
@@ 4 The NPG method with Max Line Search and Convergence Analysis
Throughout this section, let us define $l(k)\in\{k-m_{k},\ldots,k\}$ as an index satisfying
@@
Note that Assumption 3.2 (b) and Assumption 4.2 imply that $q$ is uniformly continuous on the corresponding sublevel $\mathcal{L}_{q}(x^{0})$. The requirement of uniform continuity about the objective function plays an essential role in the context of nonmonotone line search rules with the maximum taste [14]. We here recall some results for the convenience of readers.
@@
Let Assumption 3.2 and Assumption 4.2 hold and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1, then
@@
    the sequence $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$ is monotonically decreasing,
@@
    the sequences $\{x^{k}\}_{k\in\mathbb{N}}$, $\{x^{l(k)}\}_{k\in\mathbb{N}}$ $\subset\mathcal{L}_{q}(x^{0})\subset\operatorname{dom}g$,
@@
    suppose $\bar{x}$ is an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$ such that $x^{k}\to_{\mathcal{K}}\bar{x}$ holds. Then $\bar{x}$ is an M-stationary point of (Q), and $\gamma_{k}\|x^{k+1}-x^{k}\|\to_{\mathcal{K}}0$ is valid.
With the aid of Proposition 4.3 (d) as well as the techniques in Lemma 3.5, we can obtain the following result.
@@
Let Assumption 3.2 and Assumption 4.2 hold, let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1, $\bar{x}$ is an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$ and consequently let $\{x^{k}\}_{k\in\mathcal{K}}$ be a subsequence converging to some point $\bar{x}$. Then, for any $\rho>0$, there is a constant $\bar{\gamma}_{\rho}>0$ (depending on $\rho$) such that $\gamma_{k}\leq\bar{\gamma}_{\rho}$ for all $k\in\mathbb{N}$ such that $x^{k}\in B_{\rho}(\bar{x})$.
We now discuss the convergence properties of the sequence generated by Algorithm 4.1 in the presence of the KL property of $q$ at the mentioned accumulation point $\bar{x}$. Let sufficiently small $\eta>0$ be the corresponding constant from the definition of the associated desingularization function $\chi$. In view of Lemma 4.4, we can find a sufficiently large index $\hat{k}\in\mathbb{N}$ such that
@@
Consequently, let $L_{\rho}>0$ be the (global) Lipschitz constant of $\nabla f$ on $C_{\rho}$ from (4.6). In view of Lemma 4.4, let $\bar{\gamma}_{\rho}>0$ satisfy
@@
Proposition 4.5.
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1. Then there exists a constant $c_{\hat{k}}>0$ such that
$\|x^{l(k+1)}-x^{k+1}\|\leq c_{\hat{k}}\|x^{l(k+1)}-x^{l(k+1)-1}\|\quad\forall k\geq\hat{k},$
where $\hat{k}$ is defined in (4.3).
Proof.
By contradiction, for arbitrary $c>0$, there exists a $\bar{k}\geq\hat{k}$ such that
$\|x^{l(\bar{k}+1)}-x^{\bar{k}+1}\|>c\|x^{l(\bar{k}+1)}-x^{l(\bar{k}+1)-1}\|.$
Meanwhile, for the sufficiently large $\bar{k}\geq\hat{k}$, from (4.3), one has
$\|x^{l(\bar{k}+1)}-x^{\bar{k}+1}\|\leq\sum_{i=\bar{k}+1-m}^{\bar{k}+1}\|x^{i+1}-x^{i}\|\leq m\eta.$
Therefore, one has
$m\eta>c\|x^{l(\bar{k}+1)}-x^{l(\bar{k}+1)-1}\|.$
Now, choose $c:=\frac{1}{\|x^{l(\bar{k}+1)}-x^{l(\bar{k}+1)-1}\|}$, we obtain:
$m\eta>1,$
which contradicts the fact that $\eta$ is sufficiently small and $m>0$ is a fixed number. ∎
Thanks to the above notation, we can identify a new neighborhood centered at the accumulation point, with a very small radius, which is primarily used to to ensure the convergence of the whole sequence under the assumption of the KL property of $q$, we now introduce such radius.
@@
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$ with desingularization function $\chi$. Then there is a sufficiently large constant $k_{0}\in\mathcal{K}\cap\overline{K}$ such that $k_{0}-m-1$ is also sufficiently large, where $K$ is defined by (4.4), then the corresponding constant is defined as
@@
$\leq\left(\frac{8\sqrt{2}m\left(c_{\hat{k}}+1\right)(\bar{\gamma}_{\rho}+L_{\rho})}{3\delta\gamma_{\min}}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}\left(\frac{3}{m}+\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}}\right)\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\chi\big{(}q(x^{l(k_{0})})-q(\bar{x})\big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\left(\frac{4\sqrt{2}m\left(c_{\hat{k}}+1\right)}{3}\sqrt{\frac{\bar{\gamma}_{\rho}+L_{\rho}}{\mu}}+\frac{m}{3}\sqrt{\frac{2(\bar{\gamma}_{\rho}+L_{\rho})}{\delta\gamma_{\min}-\mu}}+\sqrt{\frac{\delta\gamma_{\min}}{\delta\gamma_{\min}-\mu}}\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\sqrt{\frac{2\big{(}q(x^{l(k_{0}-m-1)})-q(\bar{x})\big{)}}{\delta\gamma_{\min}}}$
@@
From the fact that $x^{k}\to_{\mathcal{K}}\bar{x}$, the decrease of $\{q(x^{l(k)})\}_{k\in\mathbb{N}}$, we have $\|x^{k_{0}}-\bar{x}\|$ is sufficiently small by the sufficiently large $k_{0}\in\mathcal{K}$, and $q(x^{l(k)})-q(\bar{x})\to 0$ from Assumption 4.2. Then we have $\chi(q(x^{l(k)})-q(\bar{x}))\to 0$ from the continuity of $\chi$ and $\chi(0)=0$. Meanwhile, because $k_{0}-m-1$ is sufficiently large, we have $q(x^{l(k_{0}-m-1)})-q(\bar{x})$ is sufficiently small. Totally, we have $\alpha<\frac{1}{2}$. ∎
Taking the similar analysis to Lemma 3.8, we obtain the following result.
@@
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, then
@@
Let Assumption 3.2 and Assumption 4.2 hold, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be any sequence generated by Algorithm 4.1. Suppose that $\{x^{k}\}_{k\in\mathcal{K}}$ is a subsequence converging to some limit point $\bar{x}$, and that $q$ has the KL property at $\bar{x}$. If $q(x^{l(k)})\neq q(\bar{x})$ for all $k\in\mathbb{N}$ and for any $v\geq k_{0}$ where $k_{0}$ is used to define $\alpha$ in Lemma 4.6, if $x^{k}\in B_{\alpha}(\bar{x})$ holds for all $k_{0}\leq k\leq v$, then we have
@@
Let $\chi:[0,\eta]\to[0,\infty)$ be the desingularization function which validates of the KL property of $q$. Due to $\chi(0)=0$ and $\chi^{\prime}(t)>0$ for all $t\in(0,\eta)$, one obtains